OpenStudy (anonymous):

The radius of a circle is 12 cm. What is the approximate circumference of the circle? (Use = 3.14.)

OpenStudy (stamp):

pi is the ratio of circumference to diameter

OpenStudy (anonymous):

these r answers 75 cm 452 cm 38 cm

OpenStudy (stamp):



OpenStudy (stamp):

Pi = Circumference / Diameter

OpenStudy (stamp):

Diameter is the length of two radius

OpenStudy (anonymous):

ok

OpenStudy (stamp):

We want the Circumference Pi = C / d C = Pi * d

OpenStudy (stamp):

diameter is two radius, d = 2r

OpenStudy (stamp):

C = Pi * d C = Pi * 2r

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

do u kno

OpenStudy (stamp):

Circumference is pi times twice the radius.

OpenStudy (stamp):

circumference of a circle radius 12 cm is http://www.wolframalpha.com/input/?i=3.14*2%2812cm%29